A normal population with unknown variance is believe to have a mean 20. Is one likely to obtain random sample of size 9 from this population that has a mean X=24 and a standard deviation of s=4.1?If not ,what conclusion would you draw?site:socratic.org


1
Expert's answer
2021-09-27T08:45:16-0400

We need to construct the "95\%" confidence interval for the population mean "\mu."

The critical value for "\alpha = 0.05" and "df = n-1 = 8" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =2.306002."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}},\bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(24-2.306\times\dfrac{4.1}{\sqrt{8}},24+2.306\times\dfrac{4.1}{\sqrt{8}})"

"=(20.848, 27.152)"

Therefore, based on the data provided, the "95\%" confidence interval for the population mean is "20.848<\mu< 27.152," which indicates that we are "95\%" confident that the true population mean "\mu" is contained by the interval "(20.848, 27.152)."

A value of "20" does not lie in the 95% confidence interval.


We need to construct the "99\%" confidence interval for the population mean "\mu."

The critical value for "\alpha = 0.01" and "df = n-1 = 8" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =3.355361."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}},\bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(24-3.355361\times\dfrac{4.1}{\sqrt{8}},24+3.355361\times\dfrac{4.1}{\sqrt{8}})"

"=(19.414, 28.587)"

Therefore, based on the data provided, the "99\%" confidence interval for the population mean is "19.414<\mu< 28.587," which indicates that we are "95\%" confident that the true population mean "\mu" is contained by the interval "(19.414, 28.587)."

A value of "20" lies in the 99% confidence interval.